OpenStudy (anonymous):

Ok so for geometry i need to answer this question but i am confuesed... Determine if triangle XYZ with coordinates X (1, 1), Y (5, 6), and Z (6, 2) is a right triangle. Use evidence to support your claim. If it is not a right triangle, what changes can be made to make it a right triangle? Be specific. Will give medal... @preetha @nincompoop

OpenStudy (amistre64):

define the distances between the points; or plot them and connect the dots

OpenStudy (amistre64):

we can determine a 90 degree angle if the slopes of connecting lines multiply to get -1 there are a few ways we can acually approach the solution.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

im thinking its more workable with the slope approach define the slope between XY = a define the slope between YZ = b define the slope between ZX = c if any 2 slopes multiply out to -1, then we have a right angle between them does ab = -1? does ac = -1? does bc = -1?

OpenStudy (anonymous):

ok im working it out now give me a couple minutes

OpenStudy (anonymous):

No none of them equal -1 but i might have done something wrong @amistre64

OpenStudy (amistre64):

what are our slopes?

OpenStudy (anonymous):

so the triangle is not a right triangle

OpenStudy (amistre64):

lets see if youve got the correct workings before we go into that :)

OpenStudy (anonymous):

xy= 5/4 yz= 1/4 zx= 1/5

OpenStudy (amistre64):

i got -4/1 not 1/4

OpenStudy (anonymous):

sorry i did zy not yz

OpenStudy (amistre64):

order of points doesnt matter